# Cosmic Ray Feedback in Massive Halos: Implications for the Distribution of Baryons

### v1

@@ 1 Introduction
The absence of ongoing star formation or a large reservoir of cool gas in galaxy groups and clusters requires a heating mechanism to balance the cooling of the hot ($T\sim 10^{7-8}~{}\mathrm{K}$) X-ray emitting plasma. Stellar feedback is energetically insufficient, so feedback from massive black holes (BHs) is generally believed to the primary heating mechanism balancing radiative losses in massive halos (Bîrzan et al., 2004; Croton et al., 2006; Fabian, 2012). In the prevailing model favored theoretically and observationally, jets from active galactic nuclei (AGN) heat the ambient halo gas. Exactly how this coupling occurs is still not fully understood, but it is likely that a combination of mechanical feedback (e.g., bubbles, turbulence, waves generated by the jet) and/or relativistic particles (cosmic rays [CRs]) created by the jet are responsible (e.g., Guo & Oh 2008; Li & Bryan 2014).
@@ 2.1 Overall Energetics
To start, we motivate why CRs from BHs should be energetically important in massive halos, particularly those with masses $M_{\rm 200}\sim 10^{12-14}M_{\odot}$. To estimate the typical central stellar and BH mass in a given halo of mass $M_{200}\gtrsim 10^{12}M_{\odot}$ (defined relative to critical) we use the $z\sim 0$ stellar mass halo mass relation from Moster et al. (2013) and the BH-stellar mass relation from Kormendy & Ho (2013). This gives $M_{\star}\simeq 10^{11}M_{\odot}(M_{200}/10^{13}M_{\odot})^{\alpha_{\star}}$ and $M_{BH}\simeq 5\times 10^{8}M_{\odot}(M_{200}/10^{13}M_{\odot})^{\alpha_{BH}}$ with $\alpha_{\star}\simeq 0.4$ and $\alpha_{BH}\simeq 0.45$. These estimates may modestly overestimate the BH mass in $\sim 10^{12}M_{\odot}$ halos because we have not accounted for the finite bulge to disk ratio; our focus, however, is on more massive halos where this is not a significant correction (Reines & Volonteri, 2015). These estimates neglect the BHs and stars not in the central galaxy, i.e., those in the rest of the group/cluster. For $\sim 10^{14}M_{\odot}$ halos, we thus likely underestimate the net stellar and BH feedback energy by a factor of $\sim 2$ (Kravtsov et al., 2018). This correction is smaller (larger) for lower (higher) halo masses.
Star formation produces CRs primarily via core-collapse supernovae [^1] Integrated over a Hubble time, Ia supernovae contribute to the total CR budget at the $\sim 10\%$ level, while fast stellar winds, colliding wind binaries, pulsars, and X-ray binaries are even smaller sources. which produce $\sim 10^{50}\epsilon_{\star,-1}$ ergs of CR energy per supernovae, where $\epsilon_{\star}=0.1\epsilon_{\star,-1}$ is the fraction of a supernova’s energy that goes into CRs. Given a typical stellar initial mass function there is roughly 1 supernova per 100 $M_{\odot}$ of star’s formed, implying that the total CR energy from supernovae over the history of star formation in a massive halo is
@@ 2.2 Spatial Distribution of Cosmic-Rays
The results that follow also neglect significant loss of CR energy from production in jets and/or quasar winds to the outer parts of halos. This approximation is most readily justified if the bulk of the CRs in massive halos are protons (as is the case in the Milky Way) because CR proton loss times are long. The composition of relativistic jets is still uncertain and they may instead be primarily electron/positron pairs (e.g., Blandford et al. 2019), which have much higher radiative and Coulomb losses. We return to this uncertainty in §5. One important distinction between diffusive and streaming transport is that in the latter case the energy in CRs is not conserved because CRs continuously heat the thermal plasma (the heating is mediated by damping of the CR-streaming-generated waves; e.g., Wentzel 1971; Zweibel 2017). We show in Appendix A, however, that the net loss of energy via this mechanism is modest in our application and so reasonable to neglect for the purposes of these initial estimates.
@@
To compare equation 3 to the thermal pressure of the halo gas, we assume that the halo baryons have the universal pressure profile of Arnaud et al. (2010), which is derived empirically as function of halo mass and radius (primarily in massive systems like galaxy clusters). Figure 3 shows the resulting ratio of the CR pressure to the thermal gas pressure at $r=R_{200}$ as a function of halo mass and CR diffusion coefficient (we assume $t\sim 10$ Gyr, consistent with most BH growth at $z\sim 1-2$). Figure 3 assumes $\epsilon_{CR,BH}=3\times 10^{-3}$ but the resulting $p_{CR}/p_{gas}$ can be scaled linearly for any other value of $\epsilon_{CR,BH}$. Note that the diffusion coefficient $\kappa$ in Figure 3 is the effective $\kappa$ in the CGM/ICM at $\sim R_{200}$, which is expected to be much larger than the Milky Way ISM value, and itself will depend on local plasma conditions and thus on halo mass (see §3).
@@ 3 CR Transport
Another way to think about CR transport via streaming in the context of halos is that the sound crossing time of a virialized halo is $\sim 0.1\,t_{\rm Hubble}$. The Alfvén crossing time is thus $\sim f_{s}^{-1}0.1\beta^{1/2}t_{\rm Hubble}$ where the effective radial speed is taken to be $f_{s}v_{A}$. Depending on the exact values of $f_{s}$ and $\beta$ (which are uncertain), CRs streaming at $v_{A}$ can thus just reach a few times the virial radius in a Hubble time. The highest diffusion coefficient model in the upper panel of Figure 4 thus requires $\beta^{1/2}f_{s}^{-1}\lesssim 10$ or that transport is in fact diffusive not purely streaming. In general there are corrections to streaming at $v_{A}$ due to the damping of the waves scattering the CRs (Skilling, 1971; Wiener et al., 2013); these corrections always increase the effective transport speed of the CRs (so $f_{s}\gtrsim 1$ is in principle possible). Even in the paradigm of scattering by waves generated by the streaming instability, transport can thus be more rapid than pure streaming at $v_{A}$. Note also that if the CR pressure builds up to become dynamically important and drive an outflow then the dominant transport becomes advection with the outflowing gas which can be faster than the Alfvén speed (as in the well-developed theory of CR-driven galactic winds, where the flow transitions from sub to super-Alfvénic as a function of radius; e.g., Ipavich 1975; Mao & Ostriker 2018; Quataert et al. 2022).
@@ 5 Summary and Discussion
It is also worth stressing that existing calibrations of outer halo feedback using galaxy group gas fractions and X-ray luminosities (e.g., van Daalen et al. 2020) do not include models in which the dominant feedback in halo outskirts is by a separate CR fluid (as we are proposing here). It is not at all clear that models with CR feedback will follow previously established correlations between the matter power spectrum at $k\sim 1\,{\rm Mpc^{-1}}$ and galaxy group gas fractions. CR feedback is fundamentally different, in that the dynamical component responsible for the feedback is long-lived and active well after the mechanical feedback produced by BHs (e.g., jets, winds, ‘bubbles’) has turned off.
@@
A natural next step in evaluating the role of CR feedback in groups is to carry out simulations of massive halos with black hole-generated CRs (as in, e.g., Wellons et al. 2023, whose zoom-in simulations included CRs from black holes in $\lesssim 10^{13}M_{\odot}$ halos). The analytic estimates presented here motivate the CR energetics and transport properties to explore in such simulations. The numerical tools for studying CR feedback in idealized and cosmological simulations exist for both diffusive and streaming transport (e.g., Butsky & Quinn 2018; Jiang & Oh 2018; Thomas & Pfrommer 2019; Hopkins et al. 2020; Ramesh et al. 2024). Such simulations would address many key questions that we can only pose here: Does CR feedback indeed produce winds at radii of order the virial radius in $10^{13-14}M_{\odot}$ halos and do such winds drive baryons to well outside the virialized halo? What is the relative role of CRs generated by the central galaxy vs. satellites (the satellites have $\sim$ twice the stellar and black hole mass of the central in $\sim 10^{14}M_{\odot}$ halos)? The same simulations will also enable quantitative observational tests of the feedback mechanism proposed here: Do models with strong CR feedback in halo outskirts predict a pion gamma-ray flux detectable with CTAs in nearby groups? Do models with strong CR feedback in group-mass systems produce galaxy clusters (which assemble from groups!) consistent with observations (a sort of ‘CR-preheating’ question)? Can CR feedback in groups indeed modify the dark matter distribution and kSZ signal at large radii while remaining consistent with X-ray constraints on the baryon fraction and gas profiles at smaller radii (as we have speculated may occur in this paper)? Does it do so equally effectively in the halos that dominate the weak lensing and kSZ signals? Fortunately the tools are in hand, guided by the estimates in this paper, to address these questions and sharpen our understanding of the impact of CRs on structure formation.
@@ Acknowledgements.
EQ thanks Alex Amon for many conversations that stimulated this work and for valuable comments on an initial draft of the paper. We also thank Philipp Kempski and Yan-Fei Jiang for useful conversations. This work was supported in part by NSF AST grant 2107872 and by Simons Investigator grants to EQ and PFH. PFH thanks the Institute for Advanced Study for sabbatical support.

### v2

@@ 1 Introduction
The absence of ongoing star formation or a large reservoir of cool gas in galaxy groups and clusters requires a heating mechanism to balance the cooling of the hot ($T\sim 10^{7-8}~{}\mathrm{K}$) X-ray emitting plasma. Stellar feedback is energetically insufficient, so feedback from massive black holes (BHs) is generally believed to be the primary heating mechanism balancing radiative losses in massive halos (Bîrzan et al., 2004; Croton et al., 2006; Fabian, 2012). In the prevailing model favored theoretically and observationally, jets from active galactic nuclei (AGN) heat the ambient halo gas. Exactly how this coupling occurs is still not fully understood, but it is likely that a combination of mechanical feedback (e.g., bubbles, turbulence, waves generated by the jet) and/or relativistic particles (cosmic rays [CRs]) created by the jet are responsible (e.g., Guo & Oh 2008; Li & Bryan 2014).
@@ 2.1 Overall Energetics
To start, we motivate why CRs from BHs should be energetically important in massive halos, particularly those with masses $M_{\rm 200}\sim 10^{12-14}M_{\odot}$. To estimate the typical central stellar and BH mass in a given halo of mass $M_{200}\gtrsim 10^{12}M_{\odot}$ (defined relative to critical) we use the $z\sim 0$ stellar mass halo mass relation from Moster et al. (2013), the bulge to total (B/T) stellar mass ratio from Fu et al. (2022), and the BH-stellar mass relation from Kormendy & Ho (2013). For massive halos with ${\rm B/T}\sim 1$, this gives $M_{\star}\simeq 10^{11}M_{\odot}(M_{200}/10^{13}M_{\odot})^{\alpha_{\star}}$ and $M_{BH}\simeq 5\times 10^{8}M_{\odot}(M_{200}/10^{13}M_{\odot})^{\alpha_{BH}}$ with $\alpha_{\star}\simeq 0.4$ and $\alpha_{BH}\simeq 0.45$. The plot in Figure 2 below includes the B/T factor which decreases the typical BH mass in $10^{12}M_{\odot}$ halos by a factor of 4 but is a negligible correction for $\gtrsim 10^{13.5}M_{\odot}$ halos. These estimates neglect the BHs and stars not in the central galaxy, i.e., those in the rest of the group/cluster. For $\sim 10^{14}M_{\odot}$ halos, we thus likely underestimate the net stellar and BH feedback energy by a factor of $\sim 2$ (Kravtsov et al., 2018). This correction is smaller (larger) for lower (higher) halo masses.
Star formation produces CRs primarily via core-collapse supernovae [^1] Integrated over a Hubble time, Ia supernovae contribute to the total CR budget at the $\sim 10\%$ level, while fast stellar winds, colliding wind binaries, pulsars, and X-ray binaries are even smaller sources. which produce $\sim 10^{50}\epsilon_{\star,-1}$ ergs of CR energy per supernovae, where $\epsilon_{\star}=0.1\epsilon_{\star,-1}$ is the fraction of a supernova’s energy that goes into CRs. Given a typical stellar initial mass function there is roughly 1 supernova per 100 $M_{\odot}$ of stars formed, implying that the total CR energy from supernovae over the history of star formation in a massive halo is
@@ 2.2 Spatial Distribution of Cosmic-Rays
The results that follow also neglect significant loss of CR energy from production in jets and/or quasar winds to the outer parts of halos. This approximation is most readily justified if the bulk of the CRs in massive halos are protons (as is the case in the Milky Way) because CR proton loss times are long. The dominant collisional/radiative loss process for $\gtrsim$ GeV protons is pion decay, for which the energy loss timescale is $\simeq 10^{10}(n/0.01\,{\rm cm^{-3}})$ yrs for an ambient gas density $n$ scaled to a typical value in groups and clusters at $\sim 0.1R_{200}$; this is much longer than the transport timescale in the models considered here. However, the composition of relativistic jets is still uncertain and they may instead be primarily electron/positron pairs (e.g., Blandford et al. 2019), which have much higher radiative and Coulomb losses. We return to this uncertainty in §5. One important distinction between diffusive and streaming transport is that in the latter case the energy in CRs is not conserved because CRs continuously heat the thermal plasma (the heating is mediated by damping of the CR-streaming-generated waves; e.g., Wentzel 1971; Zweibel 2017). We show in Appendix A, however, that the net loss of energy via this mechanism is modest in our application and so reasonable to neglect for the purposes of these initial estimates.
@@
To compare equation 3 to the thermal pressure of the halo gas, we assume that the halo baryons have the universal pressure profile of Arnaud et al. (2010), which is derived empirically as function of halo mass and radius (primarily in massive systems like galaxy clusters). Figure 3 shows the resulting ratio of the CR pressure to the thermal gas pressure at $r=R_{200}$ as a function of halo mass and CR diffusion coefficient (we assume $t\sim 10$ Gyr, consistent with most BH growth at $z\sim 1-2$). Figure 3 assumes $\epsilon_{CR,BH}=3\times 10^{-3}$ but the resulting $p_{CR}/p_{gas}$ can be scaled linearly for any other value of $\epsilon_{CR,BH}$. Note that the diffusion coefficient $\kappa$ in Figure 3 is the effective $\kappa$ in the CGM/ICM at $\sim R_{200}$, which is expected to be much larger than the Milky Way ISM value (see §3), and itself will depend on local plasma conditions and thus on halo mass.
@@ 3 CR Transport
Another way to think about CR transport via streaming in the context of halos is that the sound crossing time of a virialized halo is $\sim 0.1\,t_{\rm Hubble}$. The Alfvén crossing time is thus $\sim f_{s}^{-1}0.1\beta^{1/2}t_{\rm Hubble}$ where the effective radial speed is taken to be $f_{s}v_{A}$. Depending on the exact values of $f_{s}$ and $\beta$ (which are uncertain), CRs streaming at $v_{A}$ can thus just reach a few times the virial radius in a Hubble time. The highest diffusion coefficient model in the upper panel of Figure 4 thus requires $\beta^{1/2}f_{s}^{-1}\lesssim 10$ so that transport is in fact diffusive not purely streaming. In general there are corrections to streaming at $v_{A}$ due to the damping of the waves scattering the CRs (Skilling, 1971; Wiener et al., 2013); these corrections always increase the effective transport speed of the CRs (so $f_{s}\gtrsim 1$ is in principle possible). Even in the paradigm of scattering by waves generated by the streaming instability, transport can thus be more rapid than pure streaming at $v_{A}$. Note also that if the CR pressure builds up to become dynamically important and drive an outflow then the dominant transport becomes advection with the outflowing gas which can be faster than the Alfvén speed (as in the well-developed theory of CR-driven galactic winds, where the flow transitions from sub to super-Alfvénic as a function of radius; e.g., Ipavich 1975; Mao & Ostriker 2018; Quataert et al. 2022).
@@ 5 Summary and Discussion
It is also worth stressing that existing calibrations of outer halo feedback using galaxy group gas fractions and X-ray luminosities (e.g., van Daalen et al. 2020) do not include models in which the dominant feedback in halo outskirts is by a separate CR fluid (as we are proposing here). It is not at all clear that models with CR feedback will follow previously established correlations (van Daalen et al., 2020) between the matter power spectrum at $k\sim 1\,{\rm Mpc^{-1}}$ and galaxy group gas fractions. CR feedback is fundamentally different, in that the dynamical component responsible for the feedback is long-lived and active well after the mechanical feedback produced by BHs (e.g., jets, winds, ‘bubbles’) has turned off.
@@
For a fixed fraction of the energy produced by BHs going into CRs, simple energetics predict that the impact of BHs is likely to be the strongest in $\lesssim 10^{12}M_{\odot}$ halos (Fig. 2). Indeed, some cosmological zoom-in simulations that incorporate CR feedback find that the same dimensionless feedback parameters (e.g., analogues of our $\epsilon_{\rm BH}$; eq. 2) that can produce quenching in massive galaxies tend to overproduce feedback in lower mass halos (Byrne et al., 2024). This may imply that the efficiency of jet and CR production is itself a function of BH mass, an effect that is not included in our estimates in Figure 3. There is observational evidence that in the local Universe, radio emission from jets is dominated by the most massive BHs (Best et al., 2005). It is unclear, however, if this applies at redshifts $\sim 2-3$ where we have argued the bulk of CR energy is likely produced. Extensions of the work presented here accounting for possible BH-mass dependence of CR feedback would be useful for assessing the relative importance of SNe and BH CR feedback in $\sim 10^{12}$ halos that host MW-like galaxies. It is plausible, e.g., that BHs and SNe are similarly important for CR production in $10^{12}M_{\odot}$ halos, with BHs becoming increasingly dominant at higher halo masses.
A natural next step in evaluating the role of CR feedback in groups is to carry out simulations of massive halos with black hole-generated CRs (as in, e.g., Wellons et al. 2023; Byrne et al. 2024, whose zoom-in simulations included CRs from black holes in $\lesssim 10^{13}M_{\odot}$ halos). The analytic estimates presented here motivate the CR energetics and transport properties to explore in such simulations. The numerical tools for studying CR feedback in idealized and cosmological simulations exist for both diffusive and streaming transport (e.g., Butsky & Quinn 2018; Jiang & Oh 2018; Thomas & Pfrommer 2019; Hopkins et al. 2020; Ramesh et al. 2024). Such simulations would address many key questions that we can only pose here: Does CR feedback indeed produce winds at radii of order the virial radius in $10^{13-14}M_{\odot}$ halos and do such winds drive baryons to well outside the virialized halo? What is the relative role of CRs generated by the central galaxy vs. satellites (the satellites have $\sim$ twice the stellar and black hole mass of the central in $\sim 10^{14}M_{\odot}$ halos)? The same simulations will also enable quantitative observational tests of the feedback mechanism proposed here: Do models with strong CR feedback in halo outskirts predict a pion gamma-ray flux detectable with CTAs in nearby groups? Do models with strong CR feedback in group-mass systems produce galaxy clusters (which assemble from groups!) consistent with observations (a sort of ‘CR-preheating’ question)? Can CR feedback in groups indeed modify the dark matter distribution and kSZ signal at large radii while remaining consistent with X-ray constraints on the baryon fraction and gas profiles at smaller radii (as we have speculated may occur in this paper)? Does it do so equally effectively in the halos that dominate the weak lensing and kSZ signals? Fortunately the tools are in hand, guided by the estimates in this paper, to address these questions and sharpen our understanding of the impact of CRs on structure formation.
@@ Acknowledgements.
EQ thanks Alex Amon for many conversations that stimulated this work and for valuable comments on an initial draft of the paper. We also thank Philipp Kempski and Yan-Fei Jiang for useful conversations and the referee for useful feedback. This work was supported in part by NSF AST grant 2107872 and by Simons Investigator grants to EQ and PFH. PFH thanks the Institute for Advanced Study for sabbatical support.